OpenStudy (mtalhahassan2):

Determine the points on the graph of y=3x/x-4 where the slope of the tangent is -12/25

OpenStudy (mtalhahassan2):

Can someone help!!

OpenStudy (mtalhahassan2):

@Redneckk__Lifee

OpenStudy (mtalhahassan2):

@amoodarya

OpenStudy (welshfella):

you need to find the derivative of the function use the quotient rule

OpenStudy (mtalhahassan2):

but how where we have to put the slope

OpenStudy (welshfella):

can you do that . Have you done any calculus?

OpenStudy (welshfella):

the slope will be the derivative and you equate it to -12/25 Then solve for x

OpenStudy (mtalhahassan2):

can you plz show

OpenStudy (mtalhahassan2):

i know how to find the derivation of this equation using quotient

OpenStudy (welshfella):

y' = 3(x - 4) - 3x * 1 ------------- (x - 4)^2 = -12 ----- (x - 4)^2

OpenStudy (welshfella):

now this equlas -12/25 so we have -12 -12 ----- = ------ (x - 4)^2 25 Solve this for x.

OpenStudy (welshfella):

as the numerators are both the same we can write (x - 4)^2 = 25

OpenStudy (welshfella):

now its easy once you have found the 2 values of x plug them into the original function to find the y-values of the 2 points.

OpenStudy (welshfella):

OK?

OpenStudy (welshfella):

i have to go

OpenStudy (mtalhahassan2):

@Fools101

OpenStudy (mtalhahassan2):

@aryandecoolest

OpenStudy (mtalhahassan2):